OpenStudy (anonymous):

The triangle has an area of 40 square centimeters. find the height "h" of the triangle if the base is (2n+6) cm.

OpenStudy (anonymous):

area A= 1/2 *h*(2n+6) so, h(n+3)=40 h=40/(n+3)